# Model-Based Learning for DOA Estimation with One-Bit Single-Snapshot Sparse Arrays

### v1

@@ VI Conclusions
$\frac{\partial{\mathcal{L}}\left(\hat{\mathbf{x}},\bm{\beta}^{k}\right)}{\partial\hat{\mathbf{x}}}=$ $\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\left[\mathbf{C}(\bm{\beta}^{k})\hat{\mathbf{x}}-\mathbf{w}^{k}\right]$
@@
$+\gamma^{k}\sum_{i=1}^{N}\hat{x}_{i}\left(\left|\hat{x}_{i}\right|^{2}+\eta\right)^{\frac{\alpha}{2}-1}.$ (71)
Setting the derivative to zero, we obtain:
Since $\bm{\Lambda}(\hat{\mathbf{x}})$ is a nonlinear function of $\hat{\mathbf{x}}$, Equation (72) is inherently nonlinear, making it challenging to solve directly. However, an iterative approximation method can be utilized to efficiently estimate its solution. In this approach, at each iteration, Equation (72) is approximated as a following linear problem:
The solution of (73) is given as:
$\hat{\mathbf{x}}^{k+1}=\left[\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\mathbf{C}(\bm{\beta}^{k})+\gamma^{k}\bm{\Lambda}(\hat{\mathbf{x}}^{k})\right]^{-1}\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\mathbf{w}^{k}.$ (74)
With the estimate $\hat{\mathbf{x}}^{k+1}$ obtained from Equation (74), we can further estimate $\bm{\beta}^{k+1}$. Rearrange the objective function as:
$\mathcal{L}\left(\hat{\mathbf{x}},\bm{\beta}^{k}\right)=$ $\frac{1}{2}\bigg{(}\bm{\beta}^{\mathrm{T}}\mathrm{diag}(\hat{\mathbf{x}}^{k+1})^{*}\mathbf{\mathcal{B}}^{\mathrm{H}}\mathbf{\mathcal{B}}\mathrm{diag}(\hat{\mathbf{x}}^{k+1})\bm{\beta}$
$+(\hat{\mathbf{x}}^{k+1})^{\mathrm{H}}\mathbf{\mathcal{A}}^{\mathrm{H}}-(\mathbf{w}^{k})^{\mathrm{H}}\mathbf{\mathcal{B}}\mathrm{diag}(\hat{\mathbf{x}}^{k+1})\bm{\beta}\bigg{)}+C^{\prime},$ (75)
where $C^{\prime}$ represents the sum of all terms that are not related to $\bm{\beta}$. Taking the first-order derivative of (75) and setting it to zero, we have:
@@
$=$ $\Re\left(\mathrm{diag}(\hat{\mathbf{x}}^{k+1})^{*}\mathcal{B}^{\mathrm{H}}\left[\mathbf{w}^{k}-\mathcal{A}\hat{\mathbf{x}}^{k+1}\right]\right).$ (76)
Thus, we have the estimate of $\bm{\beta}^{k+1}$, given as:
@@
×$\times\Re\left(\mathrm{diag}(\hat{\mathbf{x}}^{k+1})^{*}\mathcal{B}^{\mathrm{H}}\left[\mathbf{w}^{k}-\mathcal{A}\hat{\mathbf{x}}^{k+1}\right]\right).$ (77)

### v2

@@ Appendix A Proof of Theorem 1
$\leq\frac{1}{\sqrt{1-\delta_{2s}}}\|\mathcal{A}\hat{\mathbf{x}}^{\dagger}-\mathbf{v}^{\dagger}\|_{2}+\left(\frac{\|\mathcal{A}\|_{2}}{\sqrt{1-\delta_{2s}}}+1\right)\sigma_{s}(\hat{\mathbf{x}}^{\dagger})_{2},$ (85)
where $\sigma_{s}(\hat{\mathbf{x}}^{\dagger})_{2}:=\inf_{\|\mathbf{y}\|_{0}\leq s}\|\hat{\mathbf{x}}^{\dagger}-\mathbf{y}\|_{2}$ denotes the best $s$-sparse approximation error. Substituting the earlier bound on the residual norm yields:
$\|\hat{\mathbf{x}}^{\dagger}-\hat{\mathbf{x}}^{\diamondsuit}\|_{2}\leq C_{1}+C_{2}\sigma_{s}(\hat{\mathbf{x}}^{\dagger})_{2},$ (86)
where $C_{1}:=\frac{1}{\sqrt{1-\delta_{2s}}}\sqrt{2(\|\hat{\mathbf{x}}^{0}\|_{\alpha}^{2}+N)}$ and $C_{2}:=\frac{\|\mathcal{A}\|_{2}}{\sqrt{1-\delta_{2s}}}+1$. This completes the proof. ∎
Appendix B Proof of Theorem 2
Proof.
For $\eta>0$ and $\gamma>0$, the boundedness of the sequence $\{\tilde{\mathbf{x}}^{k}\}_{k\geq 0}$ guarantees a subsequence $\{\tilde{\mathbf{x}}^{k_{\jmath}}\}$ converging to some limit point $\tilde{\mathbf{x}}^{\dagger}$. Since Equation (116) of Lemma 4 in the Appendix C asserts $\|\tilde{\mathbf{x}}^{k+1}-\tilde{\mathbf{x}}^{k}\|_{2}\!\to 0$, the shifted subsequence $\{\tilde{\mathbf{x}}^{k_{\jmath+1}}\}$ also converges to the same $\tilde{\mathbf{x}}^{\dagger}$.
Substituting $\tilde{\mathbf{x}}^{k_{\jmath}}$ and $\tilde{\mathbf{x}}^{k_{\jmath+1}}$ for $\tilde{\mathbf{x}}^{k}$ and $\tilde{\mathbf{x}}^{k+1}$, respectively, in the first-order derivative of the objective function in Equation (43) in the manuscript and letting $\jmath\!\to\!\infty$ yields
$\mathbf{z}\;-\;\,\mathbf{\mathcal{A}}^{\rm H}\!\bigl(\mathbf{g}(\tilde{\mathbf{x}}^{\dagger},\tilde{\mathbf{\epsilon}}^{\dagger})\bigr)=\mathbf{0},$ (87)
where ${\mathbf{z}}=[z_{1},\cdots,z_{N}]^{\mathrm{T}}$, $z_{i}=\frac{\gamma\,\tilde{x}^{\dagger}_{i}}{\bigl(\eta+|\tilde{x}^{\dagger}_{i}|^{2}\bigr)^{1-\tfrac{\alpha}{2}}},\ 1\leq i\leq N,$ and $\tilde{\mathbf{x}}^{\dagger}$ is the limit point of the objective in Equation (43).
Applying Lemma 4 in the Appendix C gives
$\bar{\mathcal{L}}_{\rm ongrid}^{\diamond}\left(\tilde{\mathbf{x}}^{\dagger},\bm{\epsilon}^{\dagger},\gamma\right)\!\;$ $\leq\;\bar{\mathcal{L}}_{\rm ongrid}^{\diamond}\left(\tilde{\mathbf{x}}^{k_{i}},\bm{\epsilon}^{\dagger},\gamma\right)\;$
$\leq\;\bar{\mathcal{L}}_{\rm ongrid}^{\diamond}\left(\tilde{\mathbf{x}}^{0},\bm{\epsilon}^{\dagger},\gamma\right)$
$\leq\;\|\tilde{\mathbf{x}}^{0}\|_{\alpha}^{2}+N.$ (88)
Consequently, it implies that
$\bigl\lVert\mathbf{g}\bigl(\tilde{\mathbf{x}}^{\dagger},\tilde{\bm{\epsilon}}^{\dagger}\bigr)\bigr\rVert_{2}$ $\leq\;\sqrt{\frac{2(a+1)}{ab^{2}}\,\bar{\mathcal{L}}_{\rm ongrid}^{\diamond}\!\bigl(\tilde{\mathbf{x}}^{\dagger},\bm{\epsilon}^{\dagger},\gamma\bigr)}$
$\leq\sqrt{\frac{2(\mathrm{a}+1)}{\mathrm{a}\mathrm{b}^{2}}\,\bigl(\lVert\tilde{\mathbf{x}}^{\diamondsuit}\rVert_{\alpha}^{2}+N\bigr)}.$ (89)
Let $S$ be the index set of non-zero entries of $\tilde{\mathbf{x}}^{\diamondsuit}$ and $S^{\dagger}$ be the set of the $s$ largest entries (in magnitude) of $\tilde{\mathbf{x}}^{\dagger}$. Since $\|\tilde{x}^{\diamondsuit}\|_{0}\leq s$ ($s$-sparse), we have
$\bigl\|\tilde{\mathbf{x}}^{\dagger}-\tilde{\mathbf{x}}^{\diamondsuit}\bigr\|_{2}\leq$ $\bigl\|(\tilde{\mathbf{x}}^{\dagger}-\tilde{\mathbf{x}}^{\diamondsuit})_{S\cup S^{\dagger}}\bigr\|_{2}+\bigl\|(\tilde{\mathbf{x}}^{\dagger})_{(S\cup S^{\dagger})^{c}}\bigr\|_{2}$
$\leq$ $\frac{1}{\sqrt{1-\delta_{2s}}}\,\sqrt{\frac{2(\mathrm{a}+1)}{\mathrm{a}\mathrm{b}^{2}}\!\bigl(\|\tilde{\mathbf{x}}^{\diamondsuit}\|_{\alpha}^{2}+N\bigr)}$
$+\Bigl(\frac{1}{\sqrt{1-\delta_{2s}}}\,\|\mathbf{\mathcal{A}}\|_{2}+1\Bigr)\sigma_{s}\!\bigl(\tilde{\mathbf{x}}^{\dagger}\bigr)_{2}$
$=$ $C_{1}^{\prime}+C_{2}^{\prime}\sigma_{s}\!\bigl(\hat{\mathbf{x}}^{\dagger}\bigr)_{2}.$ (90)
This completes the proof of Theorem 2. ∎
Appendix C Lemmas and Proofs
Lemma 1.
Let $0<\alpha\leq 1$, $\eta\geq 0$, $\gamma\geq 0$. For any $x,y\in\mathbb{C}$,
$\frac{\gamma}{\alpha}(\eta+|x|^{2})^{\!\alpha/2}-\frac{\gamma}{\alpha}(\eta+|y|^{2})^{\!\alpha/2}-\frac{\gamma\,\Re\!\big\{(x-y)\,\overline{y}\big\}}{(\eta+|x|^{2})^{1-\frac{\alpha}{2}}}\geq$
$\frac{\gamma\,|x-y|^{2}}{2(\eta+|x|^{2})^{1-\frac{\alpha}{2}}}.$ (91)
Here, $\overline{y}$ denotes the complex conjugate of $y$.
Proof.
Let $x=x_{\mathrm{R}}+jx_{\mathrm{I}}$ and $y=y_{\mathrm{R}}+jy_{\mathrm{I}}$, and define $u=[x_{\mathrm{R}},\,x_{\mathrm{I}}]^{\mathrm{T}}$, $v=[y_{\mathrm{R}},\,y_{\mathrm{I}}]^{\mathrm{T}}\in\mathbb{R}^{2}$. Then $|x|^{2}=\|u\|^{2}$, $|y|^{2}=\|v\|^{2}$, $|x-y|^{2}=\|u-v\|^{2}$, and $\Re\!\big\{(x-y)\overline{y}\big\}=(u-v)^{\mathrm{T}}v$.
Consider $g(t)=(\eta+t)^{\alpha/2}$ for $t\geq 0$. Use the arithmetic-geometric mean equality[61], we have
$(\eta+\|a\|^{2})^{1-\frac{\alpha}{2}}(\eta+\|b\|^{2})^{\frac{\alpha}{2}}\;\leq\;\Bigl(1-\frac{\alpha}{2}\Bigr)(\eta+\|a\|^{2})+\frac{\alpha}{2}(\eta+\|b\|^{2}).$ (92)
Then we compute
@@ Appendix C Lemmas and Proofs
$\frac{\gamma}{\alpha}\!\left(\eta+|x|^{2}\right)^{{\frac{\alpha}{2}}}-\frac{\gamma}{\alpha}\!\left(\eta+|y|^{2}\right)^{{\frac{\alpha}{2}}}-\frac{\gamma\,\Re\!\big\{(x-y)\,\overline{y}\big\}}{\bigl(\eta+|x|^{2}\bigr)^{\,1-\frac{\alpha}{2}}}$
$=\frac{\gamma}{\alpha}\!\left(\eta+\|u\|_{2}^{2}\right)^{\frac{\alpha}{2}}-\frac{\gamma}{\alpha}\!\left(\eta+\|v\|_{2}^{2}\right)^{\frac{\alpha}{2}}-\frac{\gamma\,(u-v)^{\mathrm{T}}v}{\bigl(\eta+|x|^{2}\bigr)^{\,1-\frac{\alpha}{2}}}$
$=\frac{\frac{\gamma}{\alpha}(\eta+\|u\|_{2}^{2})-\frac{\gamma}{\alpha}(\eta+\|v\|_{2}^{2})^{1-\frac{\alpha}{2}}(\eta+\|u\|_{2}^{2})^{\frac{\alpha}{2}}-\gamma(u-v)^{\mathrm{T}}v}{(\eta+\|u\|_{2}^{2})^{1-\frac{\alpha}{2}}}$
$\geq\frac{\frac{\gamma}{2}(\eta+\|u\|_{2}^{2})-\frac{\gamma}{2}(\eta+\|v\|_{2}^{2})-\gamma(u-v)^{\mathrm{T}}v}{(\eta+\|u\|_{2}^{2})^{1-\frac{\alpha}{2}}}$
$=\frac{\gamma\|u-v\|_{2}^{2}}{2(\eta+\|u\|_{2}^{2})^{1-\frac{\alpha}{2}}}=\frac{\gamma|x-y|^{2}}{2(\eta+|x|_{2}^{2})^{1-\frac{\alpha}{2}}}.$
Lemma 2.
Let $\hat{\mathbf{x}}^{k+1}$ be the solution from Equation (19) in the manuscript for $k=0,1,2,\cdots$. Then
$\bigl\|\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}-\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k}\bigr\|_{2}^{2}\;\leq\;2\!\left(\mathcal{L}(\hat{\mathbf{x}}^{k},\mathbf{v}^{k},\gamma)-\mathcal{L}(\hat{\mathbf{x}}^{k+1},\mathbf{v}^{k},\gamma)\right).$ (93)
Furthermore,
$\bigl\|\hat{\mathbf{x}}^{k+1}-\hat{\mathbf{x}}^{k}\bigr\|_{2}^{2}\;\leq\;C\!\left(\mathcal{L}(\hat{\mathbf{x}}^{k})-\mathcal{L}(\hat{\mathbf{x}}^{k+1})\right).$ (94)
holds for a positive constant $C$ which is bound for $\hat{\mathbf{x}}^{k},k\geq 0$.
Proof.
We first compute
$\mathcal{L}\!\bigl(\hat{\mathbf{x}}^{k},\gamma\bigr)\;-\;\mathcal{L}\!\bigl(\hat{\mathbf{x}}^{k+1},\gamma\bigr)$
$=\frac{\gamma}{\alpha}\sum_{i=1}^{N}\!\bigl(\eta+|\hat{x}^{k}_{i}|^{2}\bigr)^{\!\alpha/2}-\frac{\gamma}{\alpha}\sum_{i=1}^{N}\!\bigl(\eta+|\hat{x}^{k+1}_{i}|^{2}\bigr)^{\!\alpha/2}$
$\quad+\;\frac{1}{2}\!\Bigl(\lVert\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k}-\mathbf{v}^{k}\rVert_{2}^{2}-\lVert\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}-\mathbf{v}^{k}\rVert_{2}^{2}\Bigr)$
$=\frac{\gamma}{\alpha}\sum_{i=1}^{N}\!\bigl(\eta+|\hat{x}^{k}_{i}|^{2}\bigr)^{\!\alpha/2}-\bigl(\eta+|\hat{x}^{k+1}_{i}|^{2}\bigr)^{\!\alpha/2}$
$\quad+\;\frac{1}{2}\,\bigl\lVert\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k}-\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}\bigr\rVert_{2}^{2}$
$\quad+\Re\!\Bigl\{\bigl(\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}-\mathbf{v}^{k}\bigr)^{\mathrm{H}}\bigl(\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k}-\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}\bigr)\Bigr\}.$ (95)
Using the first-order optimality condition, we have:
$\Re\!\Bigl\{\left(\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}-\mathbf{v}^{k}\right)^{\mathrm{H}}\left(\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k}-\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}\right)\Bigr\}$
$=$ $-\sum_{i=1}^{N}\frac{\gamma\,\overline{\hat{x}_{i}^{k+1}}\bigl(\hat{x}_{i}^{k}-\hat{x}_{i}^{k+1}\bigr)}{\bigl(\eta+|\hat{x}_{i}^{k}|^{2}\bigr)^{1-\frac{\alpha}{2}}}.$ (96)
Substituting (96) into (95) and using (91) in Lemma 1 yields:
$\mathcal{L}\!\bigl(\hat{\mathbf{x}}^{k},\gamma\bigr)\;-\;\mathcal{L}\!\bigl(\hat{\mathbf{x}}^{k+1},\gamma\bigr)$
$=$ $\frac{\gamma}{\alpha}\sum_{i=1}^{N}\bigg[\!\bigl(\eta+|\hat{x}^{k}_{i}|^{2}\bigr)^{\!\alpha/2}-\bigl(\eta+|\hat{x}^{k+1}_{i}|^{2}\bigr)^{\!\alpha/2}$
$-\frac{\alpha\,\overline{\hat{x}_{i}^{k+1}}\bigl(\hat{x}_{i}^{k}-\hat{x}_{i}^{k+1}\bigr)}{\bigl(\eta+|\hat{x}_{i}^{k}|^{2}\bigr)^{1-\frac{\alpha}{2}}}\bigg]+\;\frac{1}{2}\,\bigl\lVert\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k}-\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}\bigr\rVert_{2}^{2}$
$\geq$ $\frac{1}{2}\,\bigl\lVert\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k}-\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k+1}\bigr\rVert_{2}^{2}+\sum_{i=1}^{N}\frac{\gamma\left|\hat{x}_{i}^{k}-\hat{x}_{i}^{k+1}\right|^{2}}{\bigl(\eta+|\hat{x}_{i}^{k}|^{2}\bigr)^{1-\frac{\alpha}{2}}}.$ (97)
Thus, inequality (93) holds. Since $\mathcal{L}\!\bigl(\hat{\mathbf{x}}^{k},\gamma\bigr)$ is monotonically decreasing, it thus follows that:
@@
$\left\|\hat{\mathbf{x}}^{k}\right\|_{\alpha}^{2}\leq\mathcal{L}\!\bigl(\hat{\mathbf{x}}^{k},\gamma\bigr)\leq\mathcal{L}\!\bigl(\hat{\mathbf{x}}^{0},\gamma\bigr)=\left\|\hat{\mathbf{x}}^{0}\right\|_{\alpha}^{2}$ (98)
for $k\geq 1$. Thus, there exists a positive number $\Delta$ such that $\left\|\hat{\mathbf{x}}^{k}\right\|_{\alpha}^{2}\leq\Delta$ and,
$\frac{\gamma}{\bigl(\eta+|\hat{x}_{i}^{k}|^{2}\bigr)^{1-\frac{\alpha}{2}}}\geq\frac{\gamma}{\bigl(\eta+\Delta^{2}\bigr)^{1-\frac{\alpha}{2}}}.$ (99)
Letting $\frac{1}{C}=\frac{\gamma}{\bigl(\eta+\Delta^{2}\bigr)^{1-\frac{\alpha}{2}}}$, we obtain (94) from (97). ∎
Lemma 3.
Fix the iterate $\hat{\mathbf{x}}^{k}$ from the $k$th step and define
$\mathrm{\eta}(t):=-\frac{\phi(t)}{\Phi(t)},$ (100)
where $\phi(t)$ and $\Phi(t)$ denote the PDF and CDF, respectively, of the standard normal distribution.
Let $H(u):=\int_{0}^{u}\mathrm{\eta}(t)\,dt$ and assume
$L:=\sup_{t\in\mathbb{R}}|\mathrm{\eta}^{\prime}(t)|<1.$ (101)
Define $\beta:=\tfrac{1-L}{2}>0$ and the Lyapunov function for the $v$-step:
$\Phi_{k}(\mathbf{v}):=\beta\left\|\mathbf{v}-\mathcal{A}\hat{\mathbf{x}}^{k}\right\|_{2}^{2}+\sum_{i}H(v_{i}).$ (102)
If the update is applied elementwise as
@@
$\mathbf{v}^{k}:=\overline{\mathbf{y}}\odot\left(\mathbf{D}^{k}-\mathrm{\eta}(\mathbf{D}^{k})\right),\quad\text{where }\mathbf{D}^{k}:=\overline{\mathbf{y}}\odot\mathcal{A}\hat{\mathbf{x}}^{k},$ (103)
then the Lyapunov function decreases:
$\Phi_{k}(\mathbf{v}^{k})\leq\Phi_{k}(\mathbf{v}^{k-1}).$ (104)
Consequently, for the joint Lyapunov function
$\Phi(\mathbf{x},\mathbf{v}):=$ $\beta\left\|\mathbf{v}-\mathcal{A}\hat{\mathbf{x}}^{k}\right\|_{2}^{2}+\sum_{i}H(v_{i})$
$+\frac{\gamma}{\alpha}\sum_{i=1}^{N}\left(|x_{i}^{k}|^{2}+\eta\right)^{\alpha/2},$ (105)
there will be
$\Phi(\hat{\mathbf{x}}^{k},\mathbf{v}^{k})\leq\Phi(\hat{\mathbf{x}}^{k},\mathbf{v}^{k-1}),$ (106)
Proof.
We define the coordinate-wise potential:
$\psi_{i}(v):=\beta(v-d_{i})^{2}+H(v),$ (107)
where $d_{i}^{k}:=\hat{y}_{i}\times[\mathcal{A}\hat{\mathbf{x}}^{k}]_{i}.$
Thus, $\Phi_{k}(\mathbf{v})=\sum_{i=1}^{M}\psi_{i}(v_{i})$. The derivative of $\psi_{i}(\cdot)$ is
$\psi_{i}^{\prime}(v)=2\beta(v-d_{i}^{k})+\mathrm{\eta}(v).$ (108)
Apply a Taylor expansion of $\mathrm{\eta}(v)$ around $d_{i}^{k}$:
$\mathrm{\eta}(v)\approx\mathrm{\eta}(d_{i}^{k})+\mathrm{\eta}^{\prime}(d_{i}^{k})(v-d_{i}^{k}),$ (109)
so the derivative becomes:
$\psi_{i}^{\prime}(v)\approx(2\beta+\mathrm{\eta}^{\prime}(d_{i}^{k}))(v-d_{i}^{k})+\mathrm{\eta}(d_{i}^{k}).$ (110)
Since $|\mathrm{\eta}^{\prime}(d_{i}^{k})|\leq L$ and $\beta=\tfrac{1-L}{2}$, it follows that
$\psi_{i}^{\prime\prime}(v)=2\beta+\mathrm{\eta}^{\prime}(d_{i}^{k})\geq 0\ \ (\beta>0,0\leq\mathrm{\eta}^{\prime}(d_{i}^{k})\leq 1),$ (111)
so each $\psi_{i}$ is strongly convex. The minimizer of $\psi_{i}(v)$ is characterized by the first-order condition
$\psi_{i}^{\prime}(v_{i}^{k})=(2\beta+\eta^{\prime}(d_{i}^{k}))(v_{i}^{k}-d_{i}^{k})+\eta(d_{i}^{k})=0.$ (112)
Hence,
$v_{i}^{k}=d_{i}^{k}-\tfrac{1}{2\beta+\mathrm{\eta}^{\prime}(d_{i}^{k})}\mathrm{\eta}(d_{i}^{k}),$ (113)
and choosing $v_{i}^{k}:=d_{i}^{k}-\mathrm{\eta}(d_{i}^{k})$ is a first-order approximation of the minimizer, and thus a descent step for $\psi_{i}$. By the strong convexity of $\psi_{i}(\cdot)$, we have:
$\psi_{i}(v_{i}^{k})\leq\psi_{i}(v_{i}^{k-1})-\frac{\mu}{2}(v_{i}^{k}-v_{i}^{k-1})^{2}\quad\text{for some }\mu>0.$ (114)
Summing over $i$ gives the desired inequality:
$\Phi_{k}(\mathbf{v}^{k})\leq\Phi_{k}(\mathbf{v}^{k-1}).$ (115)
Finally, since $\hat{\mathbf{x}}^{k}$ is fixed during the $\mathbf{v}$-update, the full joint Lyapunov function $\Phi(\hat{\mathbf{x}}^{k},\mathbf{v})$ decreases as well. ∎
Lemma 4.
Let $\tilde{\mathbf{x}}^{k+1}$ be the solution from Equation (46) in the manuscript for $k=0,1,2,\cdots$. Then
$\bigl\|\mathbf{\mathcal{A}}\tilde{\mathbf{x}}^{k+1}-\mathbf{\mathcal{A}}\tilde{\mathbf{x}}^{k}\bigr\|_{2}^{2}\;\leq\;\frac{2(\mathrm{a}+1)}{\mathrm{a}\mathrm{b}^{2}}\!\bigl(\bar{\mathcal{L}}_{\rm ongrid}^{\diamond}(\tilde{x}^{k})-\bar{\mathcal{L}}_{\rm ongrid}(\tilde{x}^{k+1})\bigr).$ (116)
Furthermore,
$\bigl\|\tilde{\mathbf{x}}^{k+1}-\tilde{\mathbf{x}}^{k}\bigr\|_{2}^{2}\;\leq\;C^{\prime}\!\bigl(\bar{\mathcal{L}}_{\rm ongrid}^{\diamond}(\tilde{\mathbf{x}}^{k})-\bar{\mathcal{L}}_{\rm ongrid}^{\diamond}(\tilde{\mathbf{x}}^{k+1})\bigr)$ (117)
holds for a positive constant $C^{\prime}$ which is bound for $\tilde{\mathbf{x}}^{k},k\geq 0$.
Proof.
The proof is similar to that of Lemma 2 and is omitted. ∎